### v2

# On visualisation of completions of free groups

On visualisation of completions of free groups
Yury A. Neretin [^1] The work is supported by the grant of FWF (Austrian Scientific Funds), PAT5335224.
The purpose of the note is to visualize the pronilpotent completion of a finitely generated free group as a certain subgroup in the free Lie group.
1. Introduction
1.1. Free Lie groups and pronilpotent completions of free groups. Denote by $F_{n}$ the free group with $n$ generators. Let $F_{n}^{(k)}$ be its lower central series. According a general construction of Malcev [11] (see, also its exposition in [7], Sect. 17), any finitely generated discrete nilpotent group without torsion canonically embeds as a cocompact lattice to a nilpotent Lie group. In particular, we can apply this construction to the group $F_{n}(k):=F_{n}/F_{n}^{(k)}$ and get a nilpotent Lie group, say $\mathrm{Fr}_{n}(k)$. This Lie group corresponds to the real Lie algebra $\mathfrak{fr}_{n}{(k)}$ with generators $x_{1}$, …, $x_{n}$ and relations $x_{i_{1}},[x_{i_{2}},\dots,[x_{i_{k}},x_{i_{k+1}}]\dots]]=0$ for all tuples $x_{i_{1}}$, …, $x_{i_{k+1}}$.
According Malcev [12], taking the inverse limit
$\mathrm{Fr}_{n}:=\lim\limits_{\infty\leftarrow k}\mathrm{Fr}_{n}(k),$
we get an ’infinite-dimensional Lie group’, which corresponds to the free Lie algebra with $n$ generators.
On the other hand, we have natural surjective homomorphisms $F_{n}/F_{n}^{(k+1)}\to F_{n}/F_{n}^{(k)}$. We consider the inverse limit
(1.1) $\mathcal{F}_{n}:=\lim\limits_{\infty\leftarrow k}F_{n}(k),$
of discrete nilpotent groups $F_{n}(k)=F_{n}/F_{n}^{(k)}$ and get a continual totally disconnected group $\mathcal{F}_{n}\subset\mathrm{Fr}_{n}$ (the pronilpotent completion [^2] Consider a countable discrete group $G$ without torsion. Let $G^{(k)}$ be its lower central series. Assume that $\cap_{k}G^{(k)}$ is trivial. The pronilpotent completion of $G$ is the inverse limit of discrete nilpotent groups $G/G^{(k)}$, see [1], Sect. II.4. Equivalently, consider all normal subgroups $H_{\alpha}\subset G$, such that quotient groups $G/H_{\alpha}$ are nilpotent; assume that $\cap_{\alpha}H_{\alpha}$ is trivial. If $H_{\alpha}\supset H_{\beta}$, then we have a canonical map $G/H_{\beta}\to G/H_{\alpha}$. So, we can consider the inverse limit $\lim\limits_{\leftarrow}G/H_{\alpha}$. It is the same pronilpotent completion,
We vary this definition in two ways. Considering all normal subgroups $H_{\alpha}$ such that quotient groups $G/H_{\alpha}$ are finite nilpotent groups, we get the pro-finite-nilpotent completion $\mathcal{G}^{\text{fi-ni}}$ of $G$. Considering all normal subgroups such that quotient groups are finite $p$-groups, we get the pro-$p$-completion $\mathcal{G}_{p}$ of $G$ (by definition, these groups are compact). See [17]. In fact, $\mathcal{G}^{\text{fi-ni}}$ is isomorphic to the product $\prod_{p}\mathcal{G}_{p}$ over all primes, see [17], Proposition 2.3.8.
of $F_{n}$).
A desire for transparent descriptions of such objects arises, in particular, in the context of Kohno algebras, where the question of producing explicit formulas is natural and important. Recall that Kohno algebras [8] are Lie algebras of Malcev’s Lie groups corresponding the pure braid groups. They naturally appear in numerous problems of representation theory and mathematical physics. Quite often actions of Kohno Lie algebras lift to the actions of the corresponding Lie groups, see [13]. On the other hand, Kohno algebras are normal towers of free Lie algebras (see [19]), and the corresponding Lie groups are normal towers of ’free Lie groups’ $\mathrm{Fr}_{k}$ (see [13], Theorem 3.3, Proposition 4.5).
1.2. The statement of the paper. Denote by $\mathrm{Afr}_{n}$ the algebra of formal associative (noncommutative) series in ,,variables” $\omega_{1}$, …, $\omega_{n}$. According Magnus (see, e.g., [10], Proposition I.10.1), the multiplicative group generated by $1+\omega_{j}\in\mathrm{Afr}_{n}$ is the free group $F_{n}$. Consider the tensor product $\mathrm{Afr}_{n}\otimes\mathrm{Afr}_{n}$. Consider the homomorphism of algebras (a coproduct [^3] Our $\delta$ is a coproduct in the sense of Hopf algebras, but formally we do not refer to the structure of a Hopf algebra.)
$\delta:\mathrm{Afr}_{n}\to\mathrm{Afr}_{n}\otimes\mathrm{Afr}_{n}$
defined by its values on generators
$\delta(\omega_{j})=\omega_{j}\otimes 1+1\otimes\omega_{j}+\omega_{j}\otimes\omega_{j}.$
Theorem 1.1.
a) The group $\mathrm{Fr}_{n}^{\circ}$ of elements $g\in\mathrm{Afr}_{n}$ satisfying the condition $\delta(g)=g\otimes g$ is isomorphic to $\mathrm{Fr}_{n}$.
b) The closure $\mathcal{F}_{n}$ of $F_{n}$ in $\mathrm{Fr}_{n}^{\circ}$ is the subgroup consisting of formal series with integer coefficients.
In fact, classical works by P. Hall [6] and M. Hall [5] give a realization of the pro-nilpotent completion of $F_{n}$ in terms of Malcev coordinates [11] (see Magnus, Karras, Solitar, Chap. 5 Reutenaur [16], Sect. 6.4, but the term pro-nilpotent ’completion’ in these books is not used), we identify their construction with the subgroup $\mathcal{F}_{n}$ in $\mathrm{Fr}_{n}^{\circ}$ (Corollary 3.5.b below).
We also present a similar description of the pro-$p$-completions of $F_{n}$.
1.3. The further structure of the paper. The next section contains preliminaries. In Sect. 3 we establish the main results of the paper (Theorem 3.4 and Corollary 3.5). Sect. 4 contains some additional details.
2. The free Lie group $\mathrm{Fr}_{n}$. Preliminaries
Let $G$ be a group, by $\{g,h\}$ we denote the group commutator of two elements $g$, $h\in G$:
$\{g,h\}:=g^{-1}h^{-1}gh.$
By $[x,y]$ we denote the Lie commutator, [x,y]=xyyx[x,y]=xy-yx.
2.1. The free associative algebra. Consider a metrizable commutative topological ring $\mathcal{E}$ containing rational numbers $\mathbb{Q}$}, such that the unit of $\mathbb{Q}$} is the unit of $\mathcal{E}$ (we assume that the addition and the multiplication in $\mathcal{E}$ are jointly continuous operations). Consider the subring $\mathbb{Z}$$\subset{\mathbb{Q}}\subset\mathcal{E}$ and its closure $\overline{{\mathbb{Z}}}$ in $\mathcal{E}$.
The cases, which are actually interesting for us, are:
$\mathcal{E}$ is the field of reals $\mathbb{R}$} (so $\overline{{\mathbb{Z}}}={\mathbb{Z}}$);
$\mathcal{E}$ is the field ${\mathbb{Q}}_{p}$ of $p$-adic numbers (so $\overline{{\mathbb{Z}}}$ is the ring ${\mathbb{Z}}_{p}$ of $p$-adic integers).
Denote by $\mathrm{Afr}_{n}(\mathcal{E})$ the algebra of associative (noncommutative) formal series in letters $\omega_{1}$, …, $\omega_{\alpha}$. We denote monomials by
$\omega^{\alpha}:=\omega_{\alpha_{1}}\dots\omega_{\alpha_{k}},$
so $\alpha$ ranges in the set $\mathcal{A}$ of ordered collections $(\alpha_{1},\alpha_{2},\dots,\alpha_{k})$, where $k=0$, 11, 22, $\dots$ and $1\leqslant\alpha_{j}\leqslant n$. We write elements of $\mathrm{Afr}_{n}(\mathcal{E})$ as formal series
$x=\sum_{\alpha\in\mathcal{A}}c_{\alpha}\omega^{\alpha},\qquad c_{\alpha}\in\mathcal{E}.$
Denote a degree of monomial by
$|\alpha|=|(\alpha_{1},\alpha_{2},\dots,\alpha_{k})|:=k.$
We define the lexicographic order $\prec$ on the set $\mathcal{A}$ (equivalently, on the set of monomials $\omega^{\alpha}$) assuming that the “variables” $\omega_{j}$ are ordered as $\omega_{1}\prec\omega_{2}\prec\dots\prec\omega_{n}$.
[…]
— if $|\alpha|<|\beta|$, then $\omega^{\alpha}\lessdot\omega^{\beta}$;
— if $|\alpha|=|\beta|$, then $\omega^{\alpha}\lessdot\omega^{\beta}$ is equivalent to $\omega^{\alpha}\prec\omega^{\beta}$.
Denote by $\mathrm{Afr}_{n}^{[k]}$ the subspace consisting of homogeneous elements of power $k$ and by $\mathrm{Afr}_{n}^{[>k]}$ the ideal consisting of series $\sum_{|\alpha|>k}c_{\alpha}\omega^{\alpha}$. We equip $\mathrm{Afr}_{n}$ with a topology, assuming that $\sum c_{\alpha}^{j}\omega^{\alpha}\to\sum c_{\alpha}\omega^{\alpha}$ iff we have coefficient-wise convergence $c_{\alpha}^{j}\to c_{\alpha}$ for each $\alpha$. In other words, we consider the space $\mathrm{Afr}_{n}$ as a direct product of a countable number of copies of $\mathcal{E}$ numerated by elements $\alpha\in\mathcal{A}$ and equip this space with the topology of a direct product.
Let $z_{1}$, $z_{2}\in\mathrm{Afr}_{n}(\mathcal{E})$. We write
$z_{1}=z_{2}\,\,\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>k]}\bigr{)},$
if $z_{1}-z_{2}\in\mathrm{Afr}_{n}^{[>k]}$.
For $x\in\mathrm{Afr}_{n}^{[>0]}$ and $t\in\mathcal{E}$ we have well defined
$\exp(x)$ $:=\sum_{m\geqslant 0}\frac{1}{m!}\,\,x^{m}\,\,$ $\in$   $\,\,1+\mathrm{Afr}_{n}^{[>0]};$
$\ln(1+x)$ $:=\sum_{m>0}\frac{(-1)^{m+1}}{m}\,\,x^{m}\,\,$ $\in$ $\,\,\mathrm{Afr}_{n}^{[>0]};$
$(1+x)^{t}$ $:=\sum_{m\geqslant 0}\frac{t(t-1)\dots(t-m+1)}{m!}\,\,x^{m}\,\,$ $\in$   $\,\,1+\mathrm{Afr}_{n}^{[>0]},$
actually, in each homogeneous component $\mathrm{Afr}_{n}^{[k]}$ we have a finite sum. These series satisfy the usual identities, as $\ln(\exp(x))=x$, etc.
We need the following trivial lemmas.
Lemma 2.1.
a) Let for $x=\sum_{|\alpha|>0}c_{\alpha}\omega^{\alpha}$ all coefficients $c_{\alpha}$ be integer numbers. Let $t$ also be integer. Then all coefficients of the expansion of $(1+x)^{t}$ are integer.
b) If all coefficients $c_{\alpha}$ and the exponent $t$ are contained in $\overline{{\mathbb{Z}}}$, then all coefficients of $(1+x)^{t}$ are contained in $\overline{{\mathbb{Z}}}$.
Proof. Clearly, for $t\in{\mathbb{Z}}$ the numbers
$C_{t}^{m}=\frac{t(t-1)\dots(t-m+1)}{m!}$
are integer. The map $t\mapsto C_{t}^{m}$ is continuous on $\mathcal{E}$. Since $\mathbb{Z}$} is dense in $\overline{{\mathbb{Z}}}$, we have $C_{t}^{m}\in\overline{{\mathbb{Z}}}$ for $t\in\overline{{\mathbb{Z}}}$. $\square$
Lemma 2.2.
Let $k$, $l\geqslant 1$. Let $g=1+p+\pi$, where $p\in\mathrm{Afr}_{n}^{k}$, $\pi\in\mathrm{Afr}_{n}^{[>k]}$. Let $h=1+q+\varkappa$, where $q\in\mathrm{Afr}_{n}^{l}$, $\varkappa\in\mathrm{Afr}_{n}^{[>l]}$. Then
$\{g,h\}=1+[p,q]\,\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>k+l]}\bigr{)}.$
Proof. Denote $P=p+\pi$, $Q=q+\varkappa$. Then
$(1+P)^{-1}(1+Q)^{-1}(1+P)(1+Q)=(1+P)^{-1}\bigl{(}1+(1+Q)^{-1}(1+P)(1+Q)\bigr{)}=\\ =(1+P)^{-1}\bigl{(}1+(1-Q+Q^{2}-\dots)(1+P)(1+Q)\bigr{)}.$
We have $Q^{2}P$, $QPQ$, $\dots\in\mathrm{Afr}_{n}^{[\geqslant 2l+k]}$, $P^{2}Q$, $PQP,\dots\in\mathrm{Afr}_{n}^{[\geqslant l+2k]}$ (in both cases, these elements are contained in $\mathrm{Afr}_{n}^{[>k+l]}$). So, modulo $\mathrm{Afr}_{n}^{[>k+l]}$ we get
$(1+P)^{-1}(1+P+PQ-QP)=1+PQ-QP+(-P+P^{2}-\dots)(PQ-QP)=\\ =1+PQ-QP\,\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>k+l]}\bigr{)}=1+pq-qp\,(\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>k+l]}\bigr{)}\,.\qquad\square\!\!\!\!\vphantom{.}$
2.2. The standard coproduct. We denote by $\mathrm{Afr}_{n}(\mathcal{E})\otimes\mathrm{Afr}_{n}(\mathcal{E})$ the algebra of formal series
$\sum_{\alpha,\beta\in\mathcal{A}}h_{\alpha,\beta}\,\omega^{\alpha}\otimes\omega^{\beta},\qquad h_{\alpha,\beta}\in\mathcal{E}.$
The product is defined from the condition
$\bigl{(}\omega^{\alpha}\otimes\omega^{\beta}\bigr{)}\cdot\bigl{(}\omega^{\alpha^{\prime}}\otimes\omega^{\beta^{\prime}}\bigr{)}=\omega^{\alpha}\omega^{\alpha^{\prime}}\otimes\omega^{\beta}\omega^{\beta^{\prime}}.$
The standard coproduct $\Delta$ in $\mathrm{Afr}_{n}(\mathcal{E})$ (see [16], Subsect. 1.3, [2], Definition 2.4.2) is a homomorphism of algebras over $\mathcal{E}$
$\Delta:\mathrm{Afr}_{n}(\mathcal{E})\to\mathrm{Afr}_{n}(\mathcal{E})\otimes\mathrm{Afr}_{n}(\mathcal{E})$
defined on generators by
$\Delta(\omega_{j})=\omega_{j}\otimes 1+1\otimes\omega_{j}.$
So,
$\Delta(\omega_{j}^{p})=\sum_{i=0}^{p}\frac{p!}{i!\,(p-i)!}\,\omega^{i}_{j}\otimes\omega^{p-i}_{j}.$
We say that an unshuffle of a monomial $\omega^{\alpha}$ is a splitting of the factors $\omega_{\alpha_{1}}$, …, $\omega_{\alpha_{k}}$ into two disjoint subsets, say $\alpha_{-}$, $\alpha_{+}$. In particular, we get two new monomials $\omega^{\alpha_{-}}$ and $\omega^{\alpha_{+}}$. Then
(2.1) $\Delta(\omega^{\alpha})=\sum_{(\alpha_{-},\alpha_{+})}\omega^{\alpha_{-}}\otimes\omega^{\alpha_{+}},$
where the summation is taken over all unshuffles of $\omega^{\alpha}$.
2.3. The free Lie algebra and free Lie group. Consider the free Lie algebra $\mathfrak{fr}_{n}(\mathcal{E})$ generated by $\omega_{1}$, …, $\omega_{n}$. For any multiple commutator of elements $\omega_{j}$ we assign a polynomial expression in $\omega_{1}$, …, $\omega_{n}$, assigning to any subcommutator $[x,y]$ the associative expression xyyxxy-yx. For instance,
$\omega_{1},\omega_{2}]\mapsto\omega_{1}\omega_{2}-\omega_{2}\omega_{1},\quad[[\omega_{1},\omega_{2}],\omega_{5}]\mapsto(\omega_{1}\omega_{2}-\omega_{2}\omega_{1})\omega_{5}-\omega_{5}(\omega_{1}\omega_{2}-\omega_{2}\omega_{1}),$
etc. Then we get an embedding $\mathfrak{fr}_{n}(\mathcal{E})\to\mathrm{Afr}_{n}$, see, e.g. [16], Theorem 1.4. So we consider $\mathfrak{fr}_{n}(\mathcal{E})$ as a subspace in $\mathrm{Afr}_{n}(\mathcal{E})$. Denote by $\overline{\mathfrak{fr}}_{n}(\mathcal{E})$ the closure of $\mathfrak{fr}_{n}(\mathcal{E})$ with respect to the topology of $\mathrm{Afr}_{n}(\mathcal{E})$, i.e., the Lie algebra consisting of all formal series
$z=\sum_{k>0}z_{k},\qquad\text{where $z_{k}\in\mathfrak{fr}_{n}\cap\mathrm{Afr}_{n}^{[k]}$}.$
By the Campbell–Hausdorff formula (see, e.g., [16], Chapter 3), for any $u$, $v\in\overline{\mathfrak{fr}}_{n}(\mathcal{E})$ there exists $z(u,v)\in\overline{\mathfrak{fr}}_{n}(\mathcal{E})$ such that
$\exp(u)\,\exp(v)=\exp\bigl{(}z(u,v)\bigr{)},$
so the set
$\mathrm{Fr}_{n}(\mathcal{E}):=\exp\bigl{(}\overline{\mathfrak{fr}}_{n}(\mathcal{E})\bigr{)}$
is a group with respect to the multiplication in $\mathrm{Afr}_{n}$. The following statements (K. O. Friedrichs, R. Ree) are well known, see, e.g., [16], Theorem 1.4, Theorem 3.2:
Theorem 2.3.
a) An element $z\in\mathrm{Afr}_{n}(\mathcal{E})$ is contained in $\overline{\mathfrak{fr}}_{n}(\mathcal{E})$ if and only if
$\Delta(z)=z\otimes 1+1\otimes z.$
b) An element $g\in\mathrm{Afr}_{n}(\mathcal{E})$ is contained in $\mathrm{Fr}_{n}(\mathcal{E})$ if and only if
$\Delta(g)=g\otimes g.$
Keeping in mind this theorem and (2.1), we see that $g=\sum c_{\alpha}\omega^{\alpha}\in\mathrm{Fr}_{n}(\mathcal{E})$ if and only if the coefficients $c_{\alpha}$ satisfy the system of equations
$\begin{cases}c_{\alpha}c_{\beta}=\sum_{\gamma}c_{\gamma},\end{cases}$
where $\gamma$ ranges in the set of all shuffles (see, e.g., [16]) of $\alpha$ and $\beta$. This means that we consider all monomials $\omega^{\gamma}$ equipped with splitting of factors $\omega_{\gamma_{j}}$ into two disjoint subsets, elements of the first subset form the product $\omega^{\alpha}$, elements of the second subset form the product $\omega^{\beta}$; we take sum of such $c_{\gamma}$ with multiplicities (generally, a fixed $\omega^{\gamma}$ has several admissible splittings). For instance,
$c_{1,2}\,c_{3,4}=c_{1,2,3,4}+c_{1,3,2,4}+c_{1,3,4,2}+c_{3,1,2,4}+c_{3,1,4,2}+c_{3,4,1,2}.$
Notice, that by these equations the group $\mathrm{Fr}_{n}(\mathcal{E})$ is closed in $\mathrm{Afr}_{n}(\mathcal{E})$.
2.4. A twisted coproduct. Consider the automorphism $\gamma$ of $\mathrm{Afr}_{n}(\mathcal{E})$ defined on generators by
$\gamma(\omega_{j}):=\ln(1+\omega_{j}),$
so $\gamma^{-1}(\omega_{j})=\exp(\omega_{j})-1$. The automorphism $\gamma$ preserves the filtration $\mathrm{Afr}_{n}^{[>k]}$ and is continuous. For $t\in\mathcal{E}$ we have
$\gamma\bigl{(}\exp(t\omega_{j})\bigr{)}=(1+\omega_{j})^{t}.$
Conjugating the coproduct $\Delta$ by $\gamma$, we get a homomorphism (twisted coproduct) $\delta:\mathrm{Afr}_{n}(\mathcal{E})\mapsto\mathrm{Afr}_{n}(\mathcal{E})\otimes\mathrm{Afr}_{n}(\mathcal{E})$ defined on generators by
$\delta(\omega_{j})=\omega_{j}\otimes 1+1\otimes\omega_{j}+\omega_{j}\otimes\omega_{j}.$
Indeed, we have the following chain
$\omega_{j}\xrightarrow{\gamma^{-1}}\,-1+\exp(\omega_{j})\xrightarrow{\Delta}\,-1\otimes 1+\exp(\omega_{j})\otimes\exp(\omega_{j})\,\xrightarrow{\gamma\otimes\gamma}\\ \xrightarrow{\gamma\otimes\gamma}-1\otimes 1+(1+\omega_{j})\otimes(1+\omega_{j})=\omega_{j}\otimes 1+1\otimes\omega_{j}+\omega_{j}\otimes\omega_{j}.$
Remark. We can consider an arbitrary formal series $u(x)=x+\sum_{n\geqslant 2}h_{n}x^{n}$ and an automorphism $\omega_{j}\mapsto u(\omega_{j})$, see [4], Sect. 5, and get the corresponding coproduct. Our statements (Theorem 3.4, Corollary 3.5, Corollary 3.6, Proposition 4.1, Proposition 4.2) survive if $u(x)=\ln\bigl{(}1+x+\sum_{j\geqslant 2}m_{j}x^{j}\bigr{)}$, where $m_{j}\in{\mathbb{Z}}$, but formulas below require modifications. In particular, in Theorem 3.4 and Proposition 4.1 we must write $\Xi_{L}\bigl{(}\exp(u(\omega_{1})),\dots,\exp(u(\omega_{n}))\bigr{)}$. $\boxtimes$
It is easy to see that
$\delta(\omega_{j}^{m})=\sum_{\alpha,\beta:\,0\leqslant\alpha\leqslant m,\,0\leqslant\beta\leqslant m,\,\alpha+\beta\geqslant m}\frac{m!}{(m-\alpha)!\,(m-\beta)!\,(\alpha+\beta-m)!}\,\omega_{j}^{\alpha}\otimes\omega_{j}^{\beta}.$
Consider an arbitrary monomial $\omega^{\tau}$. We color its letters green, red, and white. For each such tri-coloring $\zeta$ we denote by $\omega^{\zeta_{+}}$ the monomial obtained by omitting green letters, by $\omega^{\zeta_{-}}$ the monomial obtained by omitting red letters. Then
(2.2) $\delta(\omega^{\tau})=\sum_{\zeta}\omega^{\zeta_{+}}\otimes\omega^{\zeta_{-}},$
where the summation is taken over the set of all tri-colorings.
Notice that
$\delta(\omega^{\alpha})=\Delta(\omega^{\alpha})\,\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>|\alpha|]}\bigr{)}.$
2.5. The group $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$. Denote
$\mathrm{Fr}_{n}^{\circ}(\mathcal{E}):=\gamma(\mathrm{Fr}_{n}(\mathcal{E})).$
Clearly, we get a topological group isomorphic to $\mathrm{Fr}_{n}$. This group is the subset in $\mathrm{Afr}_{n}(\mathcal{E})$ satisfying the equation
$\delta(g)=g\otimes g.$
Keeping in mind (2.2), we observe that $g=\sum c_{\alpha}\omega^{\alpha}$ is contained in this group iff the coefficients $c_{\alpha}$ satisfy the following system of quadratic equations:
(2.3) $\begin{cases}c_{\alpha}c_{\beta}=\sum_{(\tau,\zeta)}c_{\tau},\end{cases}$
where $(\tau,\zeta)$ ranges in the set of monomials $\omega^{\tau}$ equipped with tri-colorings $\zeta$ such that $\omega^{\zeta_{+}}=\omega^{\alpha}$, $\omega^{\zeta_{-}}=\omega^{\beta}$.
Lemma 2.4.
Let $g\in\mathrm{Fr}_{n}(\mathcal{E})$, $t\in\mathcal{E}$. Then $\gamma(g)^{t}=\gamma(g^{t})$.
Proof. Denote $h:=\gamma(g)$. First, we assume $\mathcal{E}={\mathbb{R}}$. It is easy to see that for any $g\in 1+\mathrm{Afr}_{n}^{[>0]}$ roots $g^{1/m}\in 1+\mathrm{Afr}_{n}^{[>0]}$ are uniquely defined. Therefore, $\gamma$ sends $g^{1/m}$ to $h^{1/m}$. Hence the identity $\gamma(g^{t})=h^{t}$ holds for any $t\in{\mathbb{Q}}$, by continuity this is valid for all $t\in{\mathbb{R}}$.
Let us examine an arbitrary $\mathcal{E}$. Consider a homogeneous basis $\eta_{m}$ in $\mathfrak{fr}_{n}({\mathbb{Q}})$. Let $t_{m}$, $s\in\mathcal{E}$. We must verify the identity
$\exp\bigl{(}s\sum\nolimits_{m=0}^{\infty}t_{m}\eta_{m}\bigr{)}\Bigr{|}_{\omega_{j}\mapsto\ln(1+\omega_{j})}=\Bigl{(}\exp\bigl{(}\sum\nolimits_{m=0}^{\infty}t_{m}\eta_{m}\bigr{)}\Bigr{|}_{\omega_{j}\mapsto\ln(1+\omega_{j})}\Bigr{)}^{s}.$
In each homogeneous component $\mathrm{Afr}_{n}^{[k]}$ this is an equality of polynomials in $t_{m}$, $s$ with rational coefficients. This equality is valid if $\mathcal{E}={\mathbb{R}}$, and therefore it is valid for any $\mathcal{E}$. $\square$
3. The closure of the free group
3.1. Lyndon words. Consider an alphabet of letters $a_{1}$, …, $a_{n}$. We introduce the lexicographic order $\prec$ on the set of all words. If a word $u$ is a concatenation of nonempty words $u=vw$, we say that $w$ is a suffix, and $v$ is the corresponding prefix.
A Lyndon word or Lyndon–Shirshov word (see a consistent exposition in [9], Sect. 5.2) is a word that is $\prec$-less than each its suffix. Denote the set of all Lyndon words by $\mathcal{LS}_{n}$ (the abbreviation of Lyndon and Shirshov). Denote the length of $L$ by $|L|$.
Consider a Lyndon word $u$ of length >1>1 and its longest suffix $w$ that is a Lyndon word [^4] One letter is a Lyndon word, so the set of Lyndon suffixes is nonempty.. Then the corresponding prefix $v$ is also a Lyndon word (see [9], Proposition 5.1.3). We write $u=(v\circ w)$ with brackets $(\quad)$, after this we repeat the same reasoning for Lyndon words $v$, $w$, etc. So, we get a parenthesization of the word $u$ (see, e.g., [9]), something like
(3.1) $a_{1}a_{2}a_{3}a_{1}a_{3}a_{2}a_{3}\mapsto((a_{1}\circ(a_{2}\circ a_{3}))\circ((a_{1}\circ a_{3})\circ(a_{2}\circ a_{3}))).$
Now we intend to produce three different algebraic expressions from Lyndon words.
$1^{*}.$ For an arbitrary word $M$ (not only a Lyndon word), we define the monomial $\omega^{M}$ replacing all letters $a_{j}$ by the corresponding $\omega_{j}$.
$2^{*}.$ For each Lyndon word $L$ we define
$\xi_{L}(\omega_{1},\dots,\omega_{n})\in\mathfrak{fr}_{n}({\mathbb{Q}})$
by the following rule. We replace each symbol $a_{j}$ by $\omega_{j}$ and each subexpression $(p\circ q)$ in the parenthesization of $L$ by the commutator $[p,q]$. For instance, the word $a_{1}a_{2}a_{3}a_{1}a_{3}a_{2}a_{3}$ (see (3.1)) produces
$\mapsto((a_{1}\circ(a_{2}\circ a_{3}))\circ((a_{1}\circ a_{3})\circ(a_{2}\circ a_{3})))\mapsto[[\omega_{1},[\omega_{2},\omega_{3}]],[[\omega_{1},\omega_{3}],[\omega_{2},\omega_{3}]]].$
So, we get an element of the homogeneous subspace $\mathrm{Afr}_{n}^{[|L|]}$, by the construction all coefficients in this expression are integer. Then
Theorem 3.1.
Expressions $\xi_{L}(\omega_{1},\dots,\omega_{n})$ form a basis of $\mathfrak{fr}_{n}({\mathbb{Q}})$.
See, Shirshov [18], see also [9], Theorem 5.3.1. We need the following statement (it is a special case of Theorem 5.1, [16]).
Lemma 3.2.
(3.2) $\xi_{L}(\omega_{1},\dots,\omega_{n})=\omega^{L}+\sum_{M\succ L}a_{M,L}\omega^{M}.$
$3^{*}.$ Let $L$ be a Lyndon word. Let $g_{1}$, …, $g_{n}$ be elements of a group $G$. We define
$\Xi_{L}(g_{1},\dots,g_{n})$
in the following way. In $L$ we replace all letters $a_{j}$ by the corresponding $g_{j}$ and replace each subexpression $(p\circ q)$ in the parenthesization of $L$ by the group commutator $\{p,q\}$. For instance,
$\Xi_{a_{1}a_{2}a_{3}a_{1}a_{3}a_{2}a_{3}}(g_{1},g_{2},g_{3})=\{\{g_{1},\{g_{2},g_{3}\}\},\{\{g_{1},g_{3}\},\{g_{2},g_{3}\}\}\}.$
We need the following lemma, see [16], Lemma 6.10 (also, it is a straightforward corollary of Lemma 2.2 above):
Lemma 3.3.
(3.3) $\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})=1+\xi_{L}(\omega_{1},\dots,\omega_{n})\,\,\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>|L|]}\bigr{)}.$
3.2. The statements of the paper. So, we consider the multiplicative group $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})\simeq\mathrm{Fr}_{n}(\mathcal{E})$ consisting of elements $g\in\mathrm{Afr}_{n}(\mathcal{E})$ satisfying the condition $\delta(g)=g\otimes g$.
Denote by $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},{\mathbb{Z}}]$ (resp. $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$) the subgroup in $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ consisting of all $g=\sum c_{\alpha}\omega^{\alpha}$ such that $c_{\alpha}$ are contained in $\mathbb{Z}$} (resp. in $\overline{{\mathbb{Z}}}$).
Theorem 3.4.
a) Each element of the group $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ can be represented in a unique way as a Malcev type product
(3.4) $g=\prod_{L\in\mathcal{LS}_{n}}^{(\lessdot)}\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})^{t_{L}},$
[…]
b) An element $g\in\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ is contained in $\mathrm{Fr}^{\circ}_{n}[\mathcal{E},{\mathbb{Z}}]$ if and only if all $t_{L}\in{\mathbb{Z}}$.
c) An element $g\in\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ is contained in $\mathrm{Fr}^{\circ}_{n}[\mathcal{E},\overline{{\mathbb{Z}}}]$ if and only if all $t_{L}\in\overline{{\mathbb{Z}}}$.
Recall that the elements $1+\omega_{1}$, …, $1+\omega_{n}$ generate the free group $F_{n}$.
Corollary 3.5.
a) The closure of $F_{n}$ in $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ coincides with $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$.
b) The closure of $F_{n}$ in $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ consists of all products (3.4) such that $t_{L}\in\overline{{\mathbb{Z}}}$.
Theorem 3.4 is proved in the next subsection, Corollary 3.5 in Subsect. 3.
3.3. The decomposition of elements of $\mathrm{Fr}^{\circ}_{n}(\mathcal{E})$. Decompose $g\in\mathrm{Fr}^{\circ}_{n}(\mathcal{E})$ as $g=1+\sum_{k>0}g^{[k]}$, where $g^{[k]}\in\mathrm{Afr}_{n}^{[k]}$. Consider the first nonzero term, say $g^{[m]}$. We have
$\delta(1+g^{[m]}+\sum_{k>m}g^{[k]})=(1+g^{[m]}+\sum_{k>m}g^{[k]})\otimes(1+g^{[m]}+\sum_{k>m}g^{[k]}).$
Therefore,
$\delta(g^{[m]})=g^{[m]}\otimes 1+1\otimes g^{[m]}+\text{terms of degree $>m$.}$
For this reason,
$\Delta(g^{[m]})=g^{[m]}\otimes 1+1\otimes g^{[m]},$
and by Theorem 2.3 we have $g^{[m]}\in\mathfrak{fr}_{n}(\mathcal{E})$. Decompose $g^{[m]}$ by the basis $\xi_{L}$,
(3.5) $g^{[m]}=\sum_{L\in\mathcal{LS}_{n},|L|=m}t_{L}\xi_{L}(\omega_{1},\dots,\omega_{m}).$
By Lemma 3.3, we have
(3.6) $Q:=\prod_{L\in\mathcal{LS}_{n},|L|=m}\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})^{t_{L}}=\\ =1+\sum_{L\in\mathcal{LS}_{n},|L|=m}t_{L}\xi_{L}(\omega_{1},\dots,\omega_{n})\,\,\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>m]}\bigr{)}=1+g^{[m]}\,\,\Bigl{(}\mathrm{mod}\,\,\mathrm{Afr}_{n}^{[>m]}\Bigr{)}$
(by Lemma 2.4, this product is contained in $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$). Therefore,
$gQ^{-1}=1\,\,\Bigl{(}\mathrm{mod}\,\,\mathrm{Afr}_{n}^{[>m]}\Bigr{)}.$
So, we start by $m=1$ and successively kill summands of $g=1+g^{[1]}+g^{[2]}+\dots$ of degrees 1, 2, 3, ….
This proves the statement a) of Theorem 3.4.
Proofs of b), c) are identical. To be definite, we consider the case $\overline{{\mathbb{Z}}}$. By Lemma 2.1, for $t\in\overline{{\mathbb{Z}}}$ we have $\Xi_{L}(\dots)^{t}\in\mathrm{Fr}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$. We start the proof as for the statement a). All coefficients in the expression $g^{[m]}$ are contained in $\overline{{\mathbb{Z}}}$. We decompose $g^{[m]}$ by the basis $\xi_{L}=\xi_{L}(\omega_{1},\dots,\omega_{n})$, see (3.5). Let $\xi_{L}$, where $L\in\mathcal{LS}_{n}$, be the $\prec$-first summand that actually appears in the sum (3.5). By Lemma 3.2, the $\prec$-first monomial $c_{M}\omega^{M}$ that actually appears in the decomposition of $g^{(m)}$ is $t_{L}\omega^{L}$. Since $t_{L}$ is a coefficient at monomial, we have $t_{L}\in\overline{{\mathbb{Z}}}$. So,
(3.7) $g^{[m]}=t_{L}\xi_{L}(\omega_{1},\dots,\omega_{n})+\sum_{K:K\succ L,|K|=m}t_{K}\xi_{K}(\omega_{1},\dots,\omega_{n}).$
Now we pass to the element
$g^{\prime}:=g\cdot\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})^{-t_{L}}.$
Then new $(g^{\prime})^{[m]}$ is
$\sum_{K:K\succ L,|L|=m}t_{K}\xi_{K}(\omega_{1},\dots,\omega_{n})$
instead of (3.7). In the decomposition $g^{\prime}=\sum c^{\prime}_{\alpha}\omega^{\alpha}$, all coefficients $c^{\prime}_{\alpha}$ remain to be in $\overline{{\mathbb{Z}}}$. So we come to the same problem with a shorter sum (3.5). Repeating the same argument, we ’kill’ $g^{[m]}$. $\square$
3.4. The closure of the free group. By Theorem 3.4, the statements a), b) of Corollary 3.5 are equivalent. The group $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$ is closed and contains $F_{n}$.
On the other hand, for each $L\in\mathcal{LS}_{n}$ the subgroup consisting of all
$\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})^{k},\qquad\text{where $k$ ranges in ${\mathbb{Z}}$},$
is contained in $F_{n}$. Decompose this expression as
(3.8) $\sum_{m\geqslant 0}\frac{k(k-1)\dots(k-m+1)}{m!}\,\Bigl{(}\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})-1\Bigr{)}^{m}.$
The expression in big brackets has integer coefficients and is contained in $\mathrm{Afr}_{n}^{[>0]}$. Let $t\in\overline{{\mathbb{Z}}}$ and a sequence $k_{i}$ converges to $t$. Then we have coefficient-wise convergence of $\Xi_{L}(\dots)^{k_{i}}$ to $\Xi_{L}(\dots)^{t}$. Therefore, the closure of $F_{n}$ contains all elements $\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})^{t}$, where $t\in\overline{{\mathbb{Z}}}$. Hence the closure contains $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$. $\square$
3.5. Pronilpotent and pro-$p$-completions of free groups.
a) Let $\mathcal{E}={\mathbb{R}}$. Then $\mathrm{Fr}^{\circ}_{n}[\mathcal{E},\overline{{\mathbb{Z}}}]=\mathrm{Fr}^{\circ}_{n}[{\mathbb{R}},{\mathbb{Z}}]=\mathrm{Fr}^{\circ}_{n}[{\mathbb{R}},\overline{{\mathbb{Z}}}]$ is the pronilpotent completion $\mathcal{F}_{n}$ of $F_{n}$, see (1.1).
b) Let $p$ be a prime number. Let $\mathcal{E}$ be the $p$-adic field ${\mathbb{Q}}_{p}$, so $\overline{{\mathbb{Z}}}={\mathbb{Z}}_{p}$ is the ring of $p$-adic integers. As a corollary, we get the following statements (it is contained in Furusho, Enriquez [3], Lemma 6.3).
Corollary 3.6.
$\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]=\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$ is the pro-$p$-completion of $F_{n}$.
Let us verify this claim. By definition, the group $\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$ is compact. Fix positive integers $\nu$ and $m$ such that $p^{m}>\nu$. Consider the subset $U(\nu,p^{m})$ in this group consisting of series $g=1+\sum c_{\alpha}\omega^{\alpha}$ such that for $|\alpha|\leqslant\nu$ we have $c_{\alpha}\in p^{m}{\mathbb{Z}}_{p}$. It is easy to see that $U(\nu,p^{m})$ are open normal subgroups in $\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$, and these subgroups form a fundamental system of neighborhoods of the unit. So the group $\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$ is profinite, see, e.g., [17], Theorem 2.1.3.c.
Next, we notice that
$(1+x_{1}+x_{2}+\dots)^{p^{m}}=1+x_{1}^{p^{m}}+x_{2}^{p^{m}}+\dots(\operatorname{\,mod\,}p^{m}),$
Since $p^{m}>\nu$,
$\Xi(1+\omega_{1},\dots,1+\omega_{n})^{p^{m}}=1\Bigl{(}\operatorname{\,mod\,}p^{m}\mathrm{Afr}_{n}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]+\mathrm{Afr}_{n}^{[>\nu]}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]\Bigr{)}.$
This easyly implies that a product (3.4) is contained in $U(\nu,p^{m})$ if and only if $t_{L}\in p^{m}{\mathbb{Z}}_{p}$ for $|L|\leqslant\nu$ and remaining $t_{M}$ are contained in ${\mathbb{Z}}_{p}$. So, the index of our subgroup is $p^{m\sigma(\nu)}$, where $\sigma(\nu)$ is the number of Lyndon words of length $\leqslant\nu$. So $\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]/U(\nu,p^{m})$ is a $p$-group, and $\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$ is a pro-$p$-group.
Next, let $\mathfrak{F}_{n}$ be the pro-$p$-completion of $F_{n}$. By the universality of $\mathfrak{F}_{n}$, the identical map $F_{n}\to F_{n}$ induces a homomorphism $\mathfrak{F}_{n}\to\mathrm{Fr}^{\circ}_{n}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$, see [17], Proposition 3.3.6. Since $F_{n}$ is dense in $\mathrm{Fr}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$ and $\mathfrak{F}_{n}$ is compact, this homomorphism is surjective. Therefore, it is an isomorphism, see [17], Proposition 2.5.2.
c) Consider the adelic sum $\mathbb{A}$ of all $p$-adic fields ${\mathbb{Q}}_{p}$, i.e., we consider the product ${\mathbb{Q}}_{2}\times{\mathbb{Q}}_{3}\times{\mathbb{Q}}_{5}\times\dots$ and the subring $\mathbb{A}$ in this product consisting of sequences $(z_{2},z_{3},z_{5},z_{7},\dots)$ such that $|z_{p}|\leqslant 1$ for all but a finite number of primes. We consider the diagonal embedding ${\mathbb{Z}}\to\mathbb{A}$, i.e., $m\mapsto(m,m,m,\dots)$. Then $\overline{{\mathbb{Z}}}=\prod_{p}{\mathbb{Z}}_{p}$.
The group
$\mathrm{Fr}_{n}^{\circ}[\mathbb{A},\overline{{\mathbb{Z}}}]\simeq\prod_{p}\mathrm{Fr}_{n}^{\circ}[{\mathbb{Q}}_{p},{\mathbb{Z}}_{p}]$
is the pro-finite-nilpotent completion of $F_{n}$ (cf. [17], Example 2.1.6, Proposition 2.3.8).
4. Additional details
[…]
Proposition 4.1.
a) Each element of the group $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$ can be represented in a unique way as a product
(4.1) $g=\prod_{L\in\mathcal{LS}_{n}}^{{\vartriangleleft}}\Xi_{L}(1+u_{1},\dots,1+u_{n})^{t_{L}},$
[…]
b) Similar statements hold for the groups $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},{\mathbb{Z}}]$ and $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$, in these cases, $t_{L}\in{\mathbb{Z}}$ and $\overline{{\mathbb{Z}}}$ respectively.
We must explain a sense of such product. To evaluate (4.1), it is sufficient to evaluate the product modulo $\mathrm{Afr}_{n}^{[>m]}$ for each $m$. But for all Lyndon words $L$ of length >m>m, we have
$\Xi_{L}(1+u_{1},\dots,1+u_{n})^{t_{L}}=1\,\,\bigl{(}\mathrm{mod}\,\,\mathrm{Afr}_{n}^{[>m]}\bigr{)}.$
So, actually we have a product $g^{(m)}$ of finite number of factors. Clearly,
$g^{(m+1)}=g^{(m)}\,\,\bigl{(}\operatorname{\,mod\,}\,\,\mathrm{Afr}_{n}^{[>m]}\bigr{)}.$
Taking limit of $g^{(m)}$ we get our product. Clearly, it is contained in $\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$.
Proof is the same for all 3 cases.
Uniqueness. Assume that we have two representations of $g$ as a product (4.1), for a collection $t_{L}$ and a collection $t_{L}^{\prime}$. Take the largest $k$ such that for $|L|<k$ we have $t_{L}=t_{L}^{\prime}$. By Lemma 2.2,
(*) if $|M|=k$ and $|L|<k$, then $\Xi_{M}(1+\omega_{1},\dots,1+\omega_{n})$ commutes with $\Xi_{L}(1+\omega_{1},\dots,1+\omega_{n})$ modulo $\mathrm{Afr}_{n}^{[>k]}$.
Therefore, for evaluation of $g^{(k)}$ we can move the factors $\Xi_{M}(\dots)^{t_{M}}$ to the end of the product (leaving the factors $\Xi_{L}(\dots)^{t_{L}}$ at their places). Since two products (4.1) must coincide modulo $\mathrm{Afr}_{n}^{[>k]}$, we have
$\prod_{|M|=k}\Xi_{M}(\dots)^{t_{M}}=\prod_{|M|=k}\Xi_{M}(\dots)^{t^{\prime}_{M}}.$
So $t_{M}=t^{\prime}_{M}$, and we come to a contradiction.
Existence. Denote by $Q$ the set of all products (4.1). Consider $g\in\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$. Consider the largest $k$ such that there exist $h\in Q$ satisfying $g=h\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>k]}\bigr{)}$. Keeping in mind the statement (*), we observe that for any collection of $t_{M}$ the product $h\prod_{M:\,|M|=k+1}\Xi_{M}(1+\omega_{1},\dots,1+\omega_{n})^{t_{M}}$ equals to some element of $Q$ modulo $\mathrm{Afr}_{n}^{[>k+1]}$. But $g\bigl{(}\operatorname{\,mod\,}\mathrm{Afr}_{n}^{[>k+1]}\bigr{)}$ coincide with one of such products. We again come to a contradiction. $\square$
4.2. Another parametrization of pronilpotent completions.
Proposition 4.2.
a) For any collection $a_{L}\in\mathcal{E}$, where $L$ ranges in $\mathcal{LS}_{n}$, there is a unique element $g=\sum c_{\alpha}\omega^{\alpha}\in\mathrm{Fr}_{n}^{\circ}(\mathcal{E})$, such that $c_{L}=a_{L}$.
b) The statement a) holds for the groups $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},{\mathbb{Z}}]$ and $\mathrm{Fr}_{n}^{\circ}[\mathcal{E},\overline{{\mathbb{Z}}}]$. In these cases, $a_{L}\in{\mathbb{Z}}$ and $a_{L}\in\overline{{\mathbb{Z}}}$ respectively.
c) Moreover, each $c_{\alpha}$, which is not $a_{L}$, is a polynomial $P_{\alpha}(\dots)$ in $a_{M}$ with $M\lessdot\alpha$. Each exponent $t_{L}$ in the product (3.4) has the form $t_{L}=a_{L}+q_{L}$, where $q_{L}(\dots)$ is a polynomial depending on $a_{M}$ with $M\lessdot L$. All these polynomials have rational coefficients and take integer values at integer points [^6] But their coefficients are not integer..
We need the following lemma:
Lemma 4.3.
[…]
— each coefficient $c^{M}_{\alpha}$ is a polynomial $P^{M}_{\alpha}(\cdot)$ in $c_{L}$, where $L$ ranges in the set of Lyndon words $\lessdot M$;
— for all $M\gtrdot\alpha$ polynomials $P^{M}_{\alpha}$ coincide;
$t_{M}=c_{M}-P^{M}_{M}$;
— all polynomials $P^{M}_{\alpha}(\cdot)$ have rational coefficients and take integer values at integer $a_{L}$.
Proof of Lemma. The stabilization of polynomials $P^{M}_{\alpha}$ for $M\gtrdot\alpha$ follows from Lemma 3.2. We prove the remaining statements by induction. Assume that they are valid for some $g_{K}$, where $K$ is a Lyndon word. Keeping in mind Lemmas 3.3 and 3.2, we get
$g_{K_{+}}=g_{K}\cdot\Xi_{K}(1+\omega_{1},\dots,1+\omega_{n})^{t_{K}}=g_{K}\Bigl{(}1+t_{K}\omega^{K}+\sum_{\gamma\gtrdot K}\omega^{\gamma}\Bigr{)}.$
Therefore,
(4.2) $c_{K}=c^{K_{+}}_{K}=c_{K}^{K}+t_{K}.$
So we can write $g_{K_{+}}$ as
$g_{K_{+}}=g_{K}\cdot\Xi_{K}(1+\omega_{1},\dots,1+\omega_{n})^{c_{K}-c_{K}^{K}}.$
The coefficients of the second factor are polynomial expressions in $(c_{K}-P_{K}^{K}(\cdot))$. Now all statements for $M=K_{+}$ became clear.
The base of the induction is $M=\omega_{1}$ (in this case, $g_{\omega_{1}}=1$ is the empty product). $\square$
Proof of Proposition 4.2. The statement c) follows from Lemma 4.3. Uniqueness in a) also follows from c). Existence follows from (4.2), since for any $c_{K_{+}}$ we can choose $t_{K_{+}}$.
Since polynomials are integer at integer points, we have b). $\square$
Remark. There are other versions of ,,free Lie groups” [13], [14]. They are certain dense subgroups in $\mathrm{Fr}_{n}({\mathbb{R}})$ with stronger topologies, and apparently they are more interesting than $\mathrm{Fr}_{n}({\mathbb{R}})$. For such ,,free Lie group” there arise the following questions:
— for which sequences $t_{L}$ a Malcev product (3.4) is contained in this group?
— which are conditions for the coefficients $a_{L}$ from Proposition 4.2?
— to describe the closure of $F_{n}$ in this group.